OpenStudy (miss.rose):

I need help with 5 questions. I will fan and medal. @Decarr432

OpenStudy (decarr432):

Okay lets see what you got for me sweetie :D

OpenStudy (miss.rose):

Here is question one:

OpenStudy (decarr432):

What makes shapes similar? do you know :)

OpenStudy (miss.rose):

Give me a minute to solve them, then I will give you the answer I came up with and you can tell me if I am correct or not.

OpenStudy (decarr432):

What is the last integer on the answer choice 4

OpenStudy (miss.rose):

15

OpenStudy (miss.rose):

Actually, I need some help.

OpenStudy (decarr432):

Okay what do you think the answer is form first glance?

OpenStudy (decarr432):

from

OpenStudy (miss.rose):

B

OpenStudy (decarr432):

Close but no lets go over a similarity Similarity=same shape difference in length of sides by multiplication so knowing this we can look at B which I just said is incorrect

OpenStudy (decarr432):

What must you multiply 2 by to get 6? and what must you multiply 4 by to get 8?

OpenStudy (miss.rose):

3 2

OpenStudy (decarr432):

Now are they the same?

OpenStudy (miss.rose):

No.

OpenStudy (decarr432):

So are the shape similar?

OpenStudy (decarr432):

shapes*

OpenStudy (miss.rose):

No.

OpenStudy (decarr432):

Okay choose another one and lets try it

OpenStudy (miss.rose):

A

OpenStudy (decarr432):

Okay set it up like I did the first one do you remember how I did that?

OpenStudy (miss.rose):

Yes..... So what must you multiply 3 by to get 6? Answer: 2 What must you multiply 8 by to get 15? Answer: 1.875

OpenStudy (decarr432):

So does that work?

OpenStudy (decarr432):

Try B now

OpenStudy (miss.rose):

We have already tried B... I think the answer is C because: What must you multiply 3 by to get 9? Answer: 3 What must you multiply 4 by to get 12? Answer 3 Is that correct?

OpenStudy (decarr432):

My bad I meant c lol

OpenStudy (decarr432):

And yes thats the answer well done

OpenStudy (decarr432):

Review how we got the answer please so I know you understand it

OpenStudy (miss.rose):

Thank you. Here is the next question, I think the answer is B, because 4*6=24 8*3=24

OpenStudy (decarr432):

Same as before and your right haha

OpenStudy (miss.rose):

Thank you. Next Question: I need help on this one:

OpenStudy (decarr432):

Okay so what is the multiple between 3 and 12 what must you multiply to get 3 to 12

OpenStudy (miss.rose):

4

OpenStudy (decarr432):

So multiply 5 by 4 and thats your answer :D

OpenStudy (miss.rose):

20? So the answer is 20?

OpenStudy (decarr432):

Yes well done

OpenStudy (miss.rose):

Thank you. Next Question: I am going to need help on this one and the next one:

OpenStudy (miss.rose):

Opps... These are the answers:

OpenStudy (decarr432):

okay so lets see what do you think the answer is?

OpenStudy (miss.rose):

\[\frac{ 2 }{ 3 }?\]

OpenStudy (decarr432):

Why did you dome to this conclusion?

OpenStudy (miss.rose):

This was just a guess....

OpenStudy (decarr432):

come*

OpenStudy (decarr432):

Okay so think of it like this what is 24 divided by 3?

OpenStudy (miss.rose):

8

OpenStudy (decarr432):

now what is 18 divided by 3?

OpenStudy (miss.rose):

6

OpenStudy (decarr432):

Now we figured that 6*3=18 and 8*3=24 are they now the same?

OpenStudy (miss.rose):

No

OpenStudy (decarr432):

So that answer is wrong next do you think 1/6 is correct?

OpenStudy (miss.rose):

I don't know...

OpenStudy (decarr432):

lets find out what is both divided by 6?

OpenStudy (miss.rose):

24/6=4 18/6=3 They are not the same.

OpenStudy (decarr432):

I've given the answer btw XD Just think about how we are doing this

OpenStudy (miss.rose):

Is it \[\frac{ 3 }{4 }?\]

OpenStudy (decarr432):

Yes and do you know why?

OpenStudy (miss.rose):

No... Why?

OpenStudy (decarr432):

Lets break it down they are both a multiple of six correct?

OpenStudy (miss.rose):

Yes.

OpenStudy (decarr432):

How many 6's in 24? and how many in 18

OpenStudy (miss.rose):

There are 4 6's in 24. There are 3 6's in 18.

OpenStudy (decarr432):

So the figure is only 3/4's of it

OpenStudy (miss.rose):

Okay. This is the 5th and final question: I need help on this one as well:

OpenStudy (decarr432):

Okay what is the area of it now without bringing it to full scale

OpenStudy (miss.rose):

10.5cm

OpenStudy (decarr432):

Okay one second while I remember how to finish sorry been a while

OpenStudy (decarr432):

okay easier way multiply 3.5 and 3 by 120 so you get 3.5*120=? and 3*120=?

OpenStudy (miss.rose):

3.5*120=420 3*120=360

OpenStudy (decarr432):

Now how do you find the area? you multiply them together right?

OpenStudy (miss.rose):

I am not sure.... 420*360=151,200

OpenStudy (decarr432):

yep so your answer is D :)

OpenStudy (miss.rose):

THANK YOU FOR ALL YOUR HELP!!!! And I mean ALL!!!!!

OpenStudy (decarr432):

Any time you got questions i love to help

OpenStudy (miss.rose):

Thank you. Bye. I don't know your time zone, so I am just going to say have a great rest of your day, or evening. ~Miss.Rose

OpenStudy (decarr432):

Its 2:22 am lol I stayed up later to help you :D

OpenStudy (miss.rose):

Oh... Thank you. Well, I am going to let you go to bed. Good Night.

OpenStudy (decarr432):

I got work to do need more help?

OpenStudy (decarr432):

Im'a be up anyway lol